OpenStudy (anonymous):

A 0.015 kg marble moving east at 0.225 ms-1 collides head-on with a 0.030 shooter marble moving west at 0.180 ms-1. After the collision, the smaller marble moves west at 0.315 ms-1. Assume that neither marble rotates before nor after the collision and that both marbles are moving on a frictionless surface. What is the velocity of the 0.030 kg marble after the collision? What is the change in kinetic energy? What type of collision (of the 3 types) is this and WHY?

OpenStudy (jamesj):

So as ever, start with Conservation of Momentum and then take it from there.

OpenStudy (anonymous):

How do I find the change in kinetic Energy?

OpenStudy (anonymous):

help me!

OpenStudy (jamesj):

After you've found the speed of both particles before and after the collision, calculate their KEs separately and add them up. For example, the KE of the first marble before the collision is \[KE_1 = \frac{1}{2}m_1v_1^2 = \frac{1}{2} (0.015)(0.225)^2 = ... \]